OpenStudy (anonymous):

simplify 27/9 is 3/1 doesn't make sense to me?

OpenStudy (anonymous):

9 will divide into both 27 and itself. How many times does 9 go into 27?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Right, and 9 divides into itself how many times?

OpenStudy (anonymous):

\[\frac{ 27 }{ 9 }\] divide by 9 \[\frac{ 27 \div 9 }{ 9 \div 9 }\] \[\frac{ 3 }{ 1 }\]

OpenStudy (anonymous):

which is just 3

OpenStudy (anonymous):

\[\frac{ 3 }{ 1 } = 3\]

OpenStudy (anonymous):

Got it....I stopped at 3/1... Thank you

OpenStudy (anonymous):

no problem :)